OpenStudy (anonymous):

i am going crazy can sm1 tis question pleeez help me with this problem 2/3=m/12

OpenStudy (anonymous):

multiply by 12

OpenStudy (anonymous):

\[\frac{2}{3}=\frac{m}{12}\] \[12\times \frac{2}{3}=m\] \[4\times 2=m\] \[8=m\]